# Radiative decays of $P$-wave bottom baryons from light-cone sum rules

### v1

@@ I Introduction
The singly bottom baryon, consisting of a heavy $bottom$ quark and two light $up/down/strange$ quarks according to the quark model Gell-Mann:1964ewy ; Zweig:1964ruk ; Ebert:1996ec ; Gerasyuta:1999pc , is an ideal platform to study the fine structure of hadron spectra Copley:1979wj ; Karliner:2008sv ; pdg . Their searches present formidable challenges in experimental endeavors, particularly in the generation of excited states, which necessitates high beam energy and large beam luminosity. In the past two decades various experimental groups have taken great strides in the field of singly bottom baryons LHCb:2012kxf ; LHCb:2018haf ; LHCb:2018vuc ; CDF:2013pvu ; LHCb:2019soc ; LHCb:2020tqd ; LHCb:2021ssn ; LHCb:2023zpu ; CMS:2021rvl :
@@
Extensive theoretical investigations have been performed to examine the mass spectra and strong decay properties of excited bottom baryons across various phenomenological methods/models, including various quark models Garcilazo:2007eh ; Ebert:2007nw ; Roberts:2007ni ; Ortega:2012cx ; Yoshida:2015tia ; Nagahiro:2016nsx ; Wang:2018fjm ; Gutierrez-Guerrero:2019uwa ; Kawakami:2019hpp ; Xiao:2020oif ; He:2021xrh ; Wang:2017kfr , various hadronic molecular models GarciaRecio:2012db ; Liang:2014eba ; An:2017lwg ; Montana:2017kjw ; Debastiani:2017ewu ; Chen:2017xat ; Nieves:2017jjx ; Huang:2018bed ; Nieves:2019jhp ; Liang:2020dxr ; Huang:2018wgr ; Yu:2018yxl , the quark pair creation model Chen:2018orb ; Chen:2018vuc ; Yang:2018lzg ; Liang:2020hbo , the chiral perturbation theory Cheng:2006dk ; Lu:2014ina ; Cheng:2015naa , QCD sum rules Aliev:2018vye ; Aliev:2018lcs ; Wang:2020pri ; Chen:2015kpa ; Mao:2015gya ; Chen:2017sci ; Yang:2019cvw ; Yang:2020zrh ; Yang:2022oog ; Liu:2007fg ; Cui:2019dzj ; Chen:2020mpy ; Tan:2023opd , and Lattice QCD Padmanath:2013bla ; Burch:2015pka ; Padmanath:2017lng ; Can:2019wts , etc. Besides, studies on their radiative transitions are also important, especially for the bottom baryons that do not have enough phase space for strong decays. Experimentally, the BaBar and Belle Collaborations have observed three radiative decay processes Aubert:2006je ; Solovieva:2008fw ; Jessop:1998wt ; Aubert:2006rv ; Yelton:2016fqw : $\Omega_{c}^{*}\rightarrow\Omega_{c}\gamma$, $\Xi_{c}^{+}\rightarrow\Xi_{c}^{+}\gamma$, and $\Xi_{c}^{0}\rightarrow\Xi_{c}^{0}\gamma$. More experimental results from Belle-II, BESIII, and LHCb Ablikim:2019hff ; Yuan:2019zfo ; Aiola:2020yam ; Fomin:2019wuw ; Audurier:2021wqk are expected in future. In the literature Zhu:2000py ; Cheng:1992xi ; Wang:2009ic ; Tawfiq:1999cf ; Gamermann:2010ga ; Wang:2009cd ; Jiang:2015xqa ; Zhu:1998ih ; Aliev:2014bma ; Aliev:2009jt ; Aliev:2016xvq ; Aliev:2011bm ; Chow:1995nw ; Bahtiyar:2016dom ; Bahtiyar:2015sga ; Ivanov:1998wj ; Savage:1994wa ; Banuls:1999br ; Ortiz-Pacheco:2023kjn ; Bernotas:2013eia ; Cho:1994vg ; Ivanov:1999bk , theorists have employed various models to study the radiative decays of singly heavy baryons. However, most studies focused on the ground-state heavy baryons, and only a few researches Bahtiyar:2016dom ; Zhu:2000py ; Zhu:1998ih ; Wang:2009ic ; Wang:2009cd ; Aliev:2014bma ; Aliev:2009jt ; Aliev:2016xvq ; Cheng:1992xi ; Cho:1994vg ; Jiang:2015xqa ; Banuls:1999br ; Savage:1994wa ; Ivanov:1999bk ; Ivanov:1998wj ; Bernotas:2013eia ; Aliev:2011bm ; Ortiz-Pacheco:2023kjn studied the excited heavy baryons.
@@ II $P$-wave bottom baryons
$[\mathbf{\bar{3}}_{F},1,0,\lambda]$ $\Lambda_{b}$ 2.19 $0.400.26<T<$ 0.40 $1.53^{+0.14}_{-0.09}$ $\Lambda_{b}(1/2^{-})$ $5.91^{+0.17}_{-0.13}$ $4\pm 2$ $0.081^{+0.020}_{-0.018}~{}(\Lambda^{0}_{b}(1/2^{-}))$
@@
$[\mathbf{\bar{3}}_{F},0,1,\rho]$ $\Lambda_{b}$ $2.33$ $0.440.35<T<$ 0.44 $1.56^{+0.14}_{-0.16}$ $\Lambda_{b}(1/2^{-})$ $5.92^{+0.17}_{-0.19}$ $0.059^{+0.013}_{-0.013}~{}(\Lambda^{0}_{b}(1/2^{-}))$
$\Xi_{b}$ $2.29$ $T=0.45,~{}PC=34\%$ $1.72^{+0.07}_{-0.07}$ $\Xi_{b}(1/2^{-})$ $6.10^{+0.08}_{-0.08}$ $0.078^{+0.011}_{-0.009}~{}(\Xi^{-}_{b}(1/2^{-}))$
$[\mathbf{\bar{3}}_{F},1,1,\rho]$ $\Lambda_{b}$ 2.17 $0.390.26<T<$ 0.39 $1.60^{+0.11}_{-0.08}$ $\Lambda_{b}(1/2^{-})$ $5.92^{+0.13}_{-0.10}$ $8\pm 3$ $0.169^{+0.036}_{-0.027}~{}(\Lambda^{0}_{b}(1/2^{-}))$
@@
$[\mathbf{\bar{3}}_{F},2,1,\rho]$ $\Lambda_{b}$ $2.17$ $0.370.30<T<$ 0.37 $1.64^{+0.11}_{-0.11}$ $\Lambda_{b}(3/2^{-})$ $5.93^{+0.13}_{-0.13}$ $17\pm 8$ $0.136^{+0.031}_{-0.029}~{}(\Lambda^{0}_{b}(1/2^{-}))$
@@
$\Xi_{b}$ $2.31$ $0.400.31<T<$ 0.40 $1.80^{+0.09}_{-0.10}$ $\Xi_{b}(3/2^{-})$ $6.10^{+0.15}_{-0.10}$ $14\pm 7$ $0.184^{+0.038}_{-0.036}~{}(\Xi^{-}_{b}(3/2^{-}))$
$\Xi_{b}(5/2^{-})$ $6.11^{+0.15}_{-0.10}$ $0.078^{+0.160}_{-0.150}~{}(\Xi_{b}^{-}(5/2^{-}))$
$[\mathbf{6}_{F},1,0,\rho]$ $\Sigma_{b}$ 1.83 $0.340.27<T<$ 0.34 $1.31\pm 0.11$ $\Sigma_{b}(1/2^{-})$ $6.05\pm 0.12$ $3\pm 1$ $0.079\pm 0.019~{}(\Sigma^{-}_{b}(1/2^{-}))$
@@
$\Xi^{\prime}_{b}$ 1.98 $0.360.26<T<$ 0.36 $1.45\pm 0.11$ $\Xi^{\prime}_{b}(1/2^{-})$ $6.18\pm 0.12$ $3\pm 1$ $0.072\pm 0.016~{}(\Xi^{\prime-}_{b}(1/2^{-}))$
@@
$\Omega_{b}$ 2.13 $0.370.26<T<$ 0.37 $1.58\pm 0.09$ $\Omega_{b}(1/2^{-})$ $6.32\pm 0.11$ $2\pm 1$ $0.133\pm 0.028~{}(\Omega^{-}_{b}(1/2^{-}))$
@@
$[\mathbf{6}_{F},0,1,\lambda]$ $\Sigma_{b}$ $1.70$ $0.320.26<T<$ 0.32 $1.25\pm 0.10$ $\Sigma_{b}(1/2^{-})$ $6.05\pm 0.11$ $0.077\pm 0.018~{}(\Sigma^{-}_{b}(1/2^{-}))$
$\Xi^{\prime}_{b}$ $1.85$ $0.330.27<T<$ 0.33 $1.40\pm 0.09$ $\Xi^{\prime}_{b}(1/2^{-})$ $6.20\pm 0.11$ $0.069\pm 0.015~{}(\Xi^{\prime-}_{b}(1/2^{-}))$
$\Omega_{b}$ 2.00 $0.340.27<T<$ 0.34 $1.54\pm 0.09$ $\Omega_{b}(1/2^{-})$ $6.34\pm 0.11$ $0.127\pm 0.028~{}(\Omega^{-}_{b}(1/2^{-}))$
$[\mathbf{6}_{F},1,1,\lambda]$ $\Sigma_{b}$ 1.94 $0.360.29<T<$ 0.36 $1.25\pm 0.11$ $\Sigma_{b}(1/2^{-})$ $6.06\pm 0.13$ $6\pm 3$ $0.075\pm 0.016~{}(\Sigma^{-}_{b}(1/2^{-}))$
@@
$\Xi^{\prime}_{b}$ 1.97 $0.380.35<T<$ 0.38 $1.38\pm 0.09$ $\Xi^{\prime}_{b}(1/2^{-})$ $6.21\pm 0.11$ $7\pm 2$ $0.069\pm 0.012~{}(\Xi^{\prime-}_{b}(1/2^{-}))$
@@
$\Omega_{b}$ 2.00 $0.390.38<T<$ 0.39 $1.48\pm 0.07$ $\Omega_{b}(1/2^{-})$ $6.34\pm 0.10$ $6\pm 2$ $0.122\pm 0.019~{}(\Omega^{-}_{b}(1/2^{-}))$
@@
$[\mathbf{6}_{F},2,1,\lambda]$ $\Sigma_{b}$ 1.84 $0.340.27<T<$ 0.34 $1.30\pm 0.13$ $\Sigma_{b}(3/2^{-})$ $6.11\pm 0.16$ $12\pm 5$ $0.102\pm 0.028~{}(\Sigma^{-}_{b}(3/2^{-}))$
@@
$\Xi^{\prime}_{b}$ 1.96 $0.350.26<T<$ 0.35 $1.41\pm 0.12$ $\Xi^{\prime}_{b}(3/2^{-})$ $6.23\pm 0.15$ $11\pm 5$ $0.091\pm 0.023~{}(\Xi^{\prime-}_{b}(3/2^{-}))$
@@
$\Omega_{b}$ 2.08 $0.370.26<T<$ 0.37 $1.53\pm 0.10$ $\Omega_{b}(3/2^{-})$ $6.35\pm 0.13$ $10\pm 4$ $0.162\pm 0.035~{}(\Omega^{-}_{b}(3/2^{-}))$
@@ III Decays of flavor $\mathbf{\bar{3}}_{F}$ baryons
We evaluate the following two-point correlation function in order to investigate the radiative decay of the $\Xi_{b}^{0}({1/2}^{-})$ belonging to the $[\mathbf{\bar{3}}_{F},1,1,\rho]$ doublet into $\Xi_{b}^{*0}(3/2)$ and $\gamma$:
@@
$G_{\Xi_{b}^{0}[{1\over 2}^{-}]\rightarrow\Xi_{b}^{*0}\gamma}(\omega,\omega^{\prime})=g_{\Xi_{b}^{0}[{1\over 2}^{-}]\rightarrow\Xi_{b}^{*0}\gamma}\times{f_{\Xi_{b}^{0}[{1\over 2}^{-}]}f_{\Xi_{b}^{*0}}\over(\bar{\Lambda}_{\Xi_{b}^{0}[{1\over 2}^{-}]}-\omega^{\prime})(\bar{\Lambda}_{\Xi_{b}^{*0}}-\omega)}\,.$ (14)
@@
Eq. (15) and Eq. (16) contain many light-cone distribution amplitudes, whose detailed definitions and explicit expressions are meticulously documented in Refs. Ball:1998je ; Ball:2006wn ; Ball:2004rg ; Ball:1998kk ; Ball:1998sk ; Ball:1998ff ; Ball:2007rt ; Ball:2007zt . In the present study we work at the renormalization scale 1 GeV, where the QCD condensates take the following values pdg ; Yang:1993bp ; Hwang:1994vp ; Ovchinnikov:1988gk ; Narison:2002woh ; Jamin:2002ev ; Ioffe:2002be ; Shifman:2001ck ; Gimenez:2005nt :
@@
We work at the symmetric point $T_{1}=T_{2}=2T$, so $u_{0}=1/2$. Now the coupling constant $g_{\Xi_{b}^{0}[{1\over 2}^{-}]\to\Xi_{b}^{*0}\gamma}$ depends on two free parameters, the threshold value $\omega_{c}$ and the Borel mass $T$. As listed in Table 1, we choose the Borel window $0.27$ GeV $<T<0.41$ GeV with $\omega_{c}=2.32$ GeV, and plot $g_{\Xi_{b}^{0}[{1\over 2}^{-}]\to\Xi_{b}^{*0}\gamma}$ with respect to $T$ in Fig. 2, where we obtain

### v3

@@ I Introduction
The singly bottom baryon, consisting of a heavy $bottom$ quark and two light $up/down/strange$ quarks according to the quark model Gell-Mann:1964ewy ; Zweig:1964ruk ; Ebert:1996ec ; Gerasyuta:1999pc , is an ideal platform to study the fine structure of hadron spectra Copley:1979wj ; Karliner:2008sv ; pdg . Their searches present formidable challenges in experimental endeavors, particularly in the generation of excited states, which necessitates high beam energy and large beam luminosity. In the past two decades, various experimental groups have taken great strides in the field of singly bottom baryons LHCb:2012kxf ; LHCb:2018haf ; LHCb:2018vuc ; CDF:2013pvu ; LHCb:2019soc ; LHCb:2020tqd ; LHCb:2021ssn ; LHCb:2023zpu ; CMS:2021rvl :
@@
Extensive theoretical investigations have been performed to examine the mass spectra and strong decay properties of excited bottom baryons across various phenomenological methods/models, including various quark models Garcilazo:2007eh ; Ebert:2007nw ; Roberts:2007ni ; Ortega:2012cx ; Yoshida:2015tia ; Nagahiro:2016nsx ; Wang:2018fjm ; Gutierrez-Guerrero:2019uwa ; Kawakami:2019hpp ; Xiao:2020oif ; He:2021xrh ; Wang:2017kfr , various hadronic molecular models GarciaRecio:2012db ; Liang:2014eba ; An:2017lwg ; Montana:2017kjw ; Debastiani:2017ewu ; Chen:2017xat ; Nieves:2017jjx ; Huang:2018bed ; Nieves:2019jhp ; Liang:2020dxr ; Huang:2018wgr ; Yu:2018yxl , the quark pair creation model Chen:2018orb ; Chen:2018vuc ; Yang:2018lzg ; Liang:2020hbo ; Wang:2017hej , the chiral perturbation theory Cheng:2006dk ; Lu:2014ina ; Cheng:2015naa , QCD sum rules Aliev:2018vye ; Aliev:2018lcs ; Wang:2020pri ; Chen:2015kpa ; Mao:2015gya ; Chen:2017sci ; Yang:2019cvw ; Yang:2020zrh ; Yang:2022oog ; Liu:2007fg ; Cui:2019dzj ; Chen:2020mpy ; Tan:2023opd , and Lattice QCD Padmanath:2013bla ; Burch:2015pka ; Padmanath:2017lng ; Can:2019wts , etc. Besides, studies on their radiative transitions are also important, especially for the bottom baryons that do not have enough phase space for strong decays. Experimentally, the BaBar and Belle Collaborations have observed three radiative decay processes Aubert:2006je ; Solovieva:2008fw ; Jessop:1998wt ; Aubert:2006rv ; Yelton:2016fqw : $\Omega_{c}^{*}\rightarrow\Omega_{c}\gamma$, $\Xi_{c}^{+}\rightarrow\Xi_{c}^{+}\gamma$, and $\Xi_{c}^{0}\rightarrow\Xi_{c}^{0}\gamma$. More experimental results from Belle-II, BESIII, and LHCb Ablikim:2019hff ; Yuan:2019zfo ; Aiola:2020yam ; Fomin:2019wuw ; Audurier:2021wqk are expected in future. In the literature Zhu:2000py ; Cheng:1992xi ; Wang:2009ic ; Tawfiq:1999cf ; Gamermann:2010ga ; Wang:2009cd ; Jiang:2015xqa ; Zhu:1998ih ; Aliev:2014bma ; Aliev:2009jt ; Aliev:2016xvq ; Aliev:2011bm ; Chow:1995nw ; Bahtiyar:2016dom ; Bahtiyar:2015sga ; Ivanov:1998wj ; Savage:1994wa ; Banuls:1999br ; Ortiz-Pacheco:2023kjn ; Bernotas:2013eia ; Cho:1994vg ; Ivanov:1999bk , theorists have employed various models to study the radiative decays of singly heavy baryons. However, most studies focused on the ground-state heavy baryons, and only a few researches Bahtiyar:2016dom ; Zhu:2000py ; Zhu:1998ih ; Wang:2009ic ; Wang:2009cd ; Aliev:2014bma ; Aliev:2009jt ; Aliev:2016xvq ; Cheng:1992xi ; Cho:1994vg ; Jiang:2015xqa ; Banuls:1999br ; Savage:1994wa ; Ivanov:1999bk ; Ivanov:1998wj ; Bernotas:2013eia ; Aliev:2011bm ; Ortiz-Pacheco:2023kjn studied the excited heavy baryons.
@@ II $P$-wave bottom baryons
$[\mathbf{\bar{3}}_{F},1,0,\lambda]$ $\Lambda_{b}$ $2.19\pm 0.10$ $0.400.26<T<$ 0.40 $1.53^{+0.14}_{-0.09}$ $\Lambda_{b}(1/2^{-})$ $5.91^{+0.17}_{-0.13}$ $4\pm 2$ $0.081^{+0.020}_{-0.018}~{}(\Lambda^{0}_{b}(1/2^{-}))$
@@
$[\mathbf{\bar{3}}_{F},0,1,\rho]$ $\Lambda_{b}$ $2.33\pm 0.10$ $0.440.35<T<$ 0.44 $1.56^{+0.14}_{-0.16}$ $\Lambda_{b}(1/2^{-})$ $5.92^{+0.17}_{-0.19}$ $0.059^{+0.013}_{-0.013}~{}(\Lambda^{0}_{b}(1/2^{-}))$
$\Xi_{b}$ $2.29\pm 0.10$ $T=0.45/PC=34\%$ $1.72^{+0.07}_{-0.07}$ $\Xi_{b}(1/2^{-})$ $6.10^{+0.08}_{-0.08}$ $0.078^{+0.011}_{-0.009}~{}(\Xi^{-}_{b}(1/2^{-}))$
$[\mathbf{\bar{3}}_{F},1,1,\rho]$ $\Lambda_{b}$ $2.17\pm 0.10$ $0.390.26<T<$ 0.39 $1.60^{+0.11}_{-0.08}$ $\Lambda_{b}(1/2^{-})$ $5.92^{+0.13}_{-0.10}$ $8\pm 3$ $0.169^{+0.036}_{-0.027}~{}(\Lambda^{0}_{b}(1/2^{-}))$
@@
$[\mathbf{\bar{3}}_{F},2,1,\rho]$ $\Lambda_{b}$ $2.17\pm 0.10$ $0.370.30<T<$ 0.37 $1.64^{+0.11}_{-0.11}$ $\Lambda_{b}(3/2^{-})$ $5.93^{+0.13}_{-0.13}$ $17\pm 8$ $0.136^{+0.031}_{-0.029}~{}(\Lambda^{0}_{b}(1/2^{-}))$
@@
$\Xi_{b}$ $2.31\pm 0.10$ $0.400.31<T<$ 0.40 $1.80^{+0.09}_{-0.10}$ $\Xi_{b}(3/2^{-})$ $6.10^{+0.15}_{-0.10}$ $14\pm 7$ $0.184^{+0.038}_{-0.036}~{}(\Xi^{-}_{b}(3/2^{-}))$
$\Xi_{b}(5/2^{-})$ $6.11^{+0.15}_{-0.10}$ $0.078^{+0.016}_{-0.015}~{}(\Xi_{b}^{-}(5/2^{-}))$
$[\mathbf{6}_{F},1,0,\rho]$ $\Sigma_{b}$ $1.83\pm 0.10$ $0.340.27<T<$ 0.34 $1.31\pm 0.11$ $\Sigma_{b}(1/2^{-})$ $6.05\pm 0.12$ $3\pm 1$ $0.079\pm 0.019~{}(\Sigma^{-}_{b}(1/2^{-}))$
@@
$\Xi^{\prime}_{b}$ $1.98\pm 0.10$ $0.360.26<T<$ 0.36 $1.45\pm 0.11$ $\Xi^{\prime}_{b}(1/2^{-})$ $6.18\pm 0.12$ $3\pm 1$ $0.072\pm 0.016~{}(\Xi^{\prime-}_{b}(1/2^{-}))$
@@
$\Omega_{b}$ $2.13\pm 0.10$ $0.370.26<T<$ 0.37 $1.58\pm 0.09$ $\Omega_{b}(1/2^{-})$ $6.32\pm 0.11$ $2\pm 1$ $0.133\pm 0.028~{}(\Omega^{-}_{b}(1/2^{-}))$
@@
$[\mathbf{6}_{F},0,1,\lambda]$ $\Sigma_{b}$ $1.70\pm 0.10$ $0.320.26<T<$ 0.32 $1.25\pm 0.10$ $\Sigma_{b}(1/2^{-})$ $6.05\pm 0.11$ $0.077\pm 0.018~{}(\Sigma^{-}_{b}(1/2^{-}))$
$\Xi^{\prime}_{b}$ $1.85\pm 0.10$ $0.330.27<T<$ 0.33 $1.40\pm 0.09$ $\Xi^{\prime}_{b}(1/2^{-})$ $6.20\pm 0.11$ $0.069\pm 0.015~{}(\Xi^{\prime-}_{b}(1/2^{-}))$
$\Omega_{b}$ $2.00\pm 0.10$ $0.340.27<T<$ 0.34 $1.54\pm 0.09$ $\Omega_{b}(1/2^{-})$ $6.34\pm 0.11$ $0.127\pm 0.028~{}(\Omega^{-}_{b}(1/2^{-}))$
$[\mathbf{6}_{F},1,1,\lambda]$ $\Sigma_{b}$ $1.94\pm 0.10$ $0.360.29<T<$ 0.36 $1.25\pm 0.11$ $\Sigma_{b}(1/2^{-})$ $6.06\pm 0.13$ $6\pm 3$ $0.075\pm 0.016~{}(\Sigma^{-}_{b}(1/2^{-}))$
@@
$\Xi^{\prime}_{b}$ $1.97\pm 0.10$ $0.380.35<T<$ 0.38 $1.38\pm 0.09$ $\Xi^{\prime}_{b}(1/2^{-})$ $6.21\pm 0.11$ $7\pm 2$ $0.069\pm 0.012~{}(\Xi^{\prime-}_{b}(1/2^{-}))$
@@
$\Omega_{b}$ $2.00\pm 0.10$ $0.390.38<T<$ 0.39 $1.48\pm 0.07$ $\Omega_{b}(1/2^{-})$ $6.34\pm 0.10$ $6\pm 2$ $0.122\pm 0.019~{}(\Omega^{-}_{b}(1/2^{-}))$
@@
$[\mathbf{6}_{F},2,1,\lambda]$ $\Sigma_{b}$ $1.84\pm 0.10$ $0.340.27<T<$ 0.34 $1.30\pm 0.13$ $\Sigma_{b}(3/2^{-})$ $6.11\pm 0.16$ $12\pm 5$ $0.102\pm 0.028~{}(\Sigma^{-}_{b}(3/2^{-}))$
@@
$\Xi^{\prime}_{b}$ $1.96\pm 0.10$ $0.350.26<T<$ 0.35 $1.41\pm 0.12$ $\Xi^{\prime}_{b}(3/2^{-})$ $6.23\pm 0.15$ $11\pm 5$ $0.091\pm 0.023~{}(\Xi^{\prime-}_{b}(3/2^{-}))$
@@
$\Omega_{b}$ $2.08\pm 0.10$ $0.370.26<T<$ 0.37 $1.53\pm 0.10$ $\Omega_{b}(3/2^{-})$ $6.35\pm 0.13$ $10\pm 4$ $0.162\pm 0.035~{}(\Omega^{-}_{b}(3/2^{-}))$
@@ III Decays of flavor $\mathbf{\bar{3}}_{F}$ baryons
We evaluate the following two-point correlation function in order to investigate the radiative decay of the $\Xi_{b}^{0}({1/2}^{-})$ belonging to the $[\mathbf{\bar{3}}_{F},1,1,\rho]$ doublet into $\Xi_{b}^{*0}(3/2^{+})$ and $\gamma$:
@@
$G_{\Xi_{b}^{0}[{1\over 2}^{-}]\rightarrow\Xi_{b}^{*0}\gamma}(\omega,\omega^{\prime})=g_{\Xi_{b}^{0}[{1\over 2}^{-}]\rightarrow\Xi_{b}^{*0}\gamma}\times{f_{\Xi_{b}^{0}[{1\over 2}^{-}]}f_{\Xi_{b}^{*0}}\over(\bar{\Lambda}_{\Xi_{b}^{0}[{1\over 2}^{-}]}-\omega^{\prime})(\bar{\Lambda}_{\Xi_{b}^{*0}}-\omega)}+\cdots\,,$ (14)
where $\cdots$ represents the contributions from excited states and the continuum.
@@
Figure 2: Feynman diagrams for the process $\Xi_{b}^{0}({1/2}^{-})\to\Xi_{b}^{*0}(3/2^{+})\gamma$.
Eq. (15) and Eq. (16) contain many light-cone distribution amplitudes, whose detailed definitions and explicit expressions are meticulously documented in Refs. Ball:1998je ; Ball:2006wn ; Ball:2004rg ; Ball:1998kk ; Ball:1998sk ; Ball:1998ff ; Ball:2007rt ; Ball:2007zt . For completeness, we also list them in Appendix A. In the present study we work at the renormalization scale 1 GeV and take into account the light-cone photon distribution amplitudes up to twist-4. We note that Eq. (15) and Eq. (16) only contain twist-2 distribution amplitudes ($\phi_{\gamma}$) and twist-3 distribution amplitudes ($\psi^{\alpha}$, $\mathcal{A}$, and $\mathcal{V}$), while they do not contain twist-4 distribution amplitudes ($h_{\gamma}$, $S$, $\widetilde{S}$, and $T_{1,2,3,4}$). As illustrated in Fig. 2, there are a total of eight Feynman diagrams, including the perturbative term, the quark condensate, the quark-gluon mixed condensate, and their combinations. We have taken into account the up and strange quark contributions to the photon DA. The QCD condensates take the following values pdg ; Yang:1993bp ; Hwang:1994vp ; Ovchinnikov:1988gk ; Narison:2002woh ; Jamin:2002ev ; Ioffe:2002be ; Shifman:2001ck ; Gimenez:2005nt :
@@
As defined in Eq. (11), $\omega$ and $\omega^{\prime}$ are momenta of the final and initial baryons, respectively. These two momenta are comparable to each other, allowing us to work at the symmetric point $T_{1}=T_{2}=2T$, so that $u_{0}=1/2$. Now the coupling constant $g_{\Xi_{b}^{0}[{1\over 2}^{-}]\to\Xi_{b}^{*0}\gamma}$ depends on two free parameters, the threshold value $\omega_{c}$ and the Borel mass $T$. As listed in Table 1, we choose the Borel window $0.27$ GeV $<T<0.41$ GeV with $\omega_{c}=2.32$ GeV, and plot $g_{\Xi_{b}^{0}[{1\over 2}^{-}]\to\Xi_{b}^{*0}\gamma}$ with respect to $T$ in Fig. 3, where we obtain